### v1

@@ 4 Extensions
4.1 Nonlinear estimand $\tau(\theta)$
@@
This subsection extends the framework to smooth, potentially nonlinear $\tau(\theta)$, and consider a local asymptotic framework in the spirit of Andrews et al. (2020); Bonhomme and Weidner (2022).
Setup.
Take any given $(\mathcal{E},\Sigma)$ and $\gamma$. Let $\tilde{\theta}^{\mathrm{obs}}$ and $\tilde{\theta}^{\mathrm{exp}}$ denote the corresponding observational and experimental estimators of $\theta\in\mathbb{R}^{p}$. Assume a common $\sqrt{n}$ rate and the following local asymptotics:
$\sqrt{n}\!\left(\tilde{\theta}^{\mathrm{exp}}-\theta\right)$ $\;\overset{d}{\longrightarrow}\;\mathcal{N}(0,\Sigma^{\mathrm{exp}}),$
$\sqrt{n}\!\left(\tilde{\theta}^{\mathrm{obs}}-\theta-b_{n}\right)$ $\;\overset{d}{\longrightarrow}\;\mathcal{N}(0,\Sigma^{\mathrm{obs}}),\qquad b_{n}\to 0,$
where $b_{n}\in\mathbb{R}^{p}$ is an unknown drift capturing local misspecification, and $\Sigma^{\mathrm{exp}},\Sigma^{\mathrm{obs}}$ are finite positive–definite matrices. We take the experimental and observational estimators to be asymptotically independent with common rates; this is not necessary but simplifies exposition. Thus $\tilde{\theta}^{\mathrm{obs}}\stackrel{{\scriptstyle p}}{{\to}}\theta$ (consistency), while $\sqrt{n}\,b_{n}$ may be bounded or even diverge. To justify first order linearization when $\sqrt{n}\,b_{n}$ may grow, we impose $\sqrt{n}\,\|b_{n}\|^{2}\to 0$ (e.g., $\|b_{n}\|=n^{-\alpha}$ with $\alpha\in(1/4,1/2]$). This implies that even if $b_{n}$ converges to zero, the asymptotic bias is possibly comparable to, smaller or larger than the asymptotic variance, therefore imposing weak restrictions on the magnitude of $\sqrt{n}b_{n}$ relative to the asymptotic variance.
Linearization of $\tau$.
@@
Let $\tau:\mathbb{R}^{p}\to\mathbb{R}$ be twice continuously differentiable in a neighborhood of $\theta$, with Lipschitz gradient and bounded Hessian. A Taylor expansion around $\theta$ yields
$\tau(\hat{\theta}_{\gamma})-\tau(\theta)=\omega(\theta)^{\top}(\hat{\theta}_{\gamma}-\theta)\;+\;\tfrac{1}{2}(\hat{\theta}_{\gamma}-\theta)^{\top}H_{\tau}(\bar{\theta})\,(\hat{\theta}_{\gamma}-\theta),$
[…]
$\sqrt{n}\bigl(\tau(\hat{\theta}_{\gamma})-\tau(\theta)\bigr)=\omega(\theta)^{\top}\sqrt{n}\,(\hat{\theta}_{\gamma}-\theta)\;+\;o_{p}(1).$ (7)
Using the fact that $\tilde{\theta}^{\mathrm{obs}}\rightarrow_{p}\theta$, we can consistently estimate the asymptotic MSE $n\mathbb{E}\left[(\tau(\hat{\theta}_{\gamma})-\tau(\theta))^{2}\right]$ by replacing $\omega(\theta)$ with $\omega(\tilde{\theta}^{\mathrm{obs}})\rightarrow_{p}\omega(\theta)$ and ignoring the $o_{p}(1)$ on the right-hand side. As a result all our results continue to hold under such linearization up-to a negligible $o_{p}(1)$.
@@ 4.2 Confidence interval length as a regret criterion
In this subsection we consider an alternative loss based on the length of confidence intervals (CIs). For a candidate design $(\mathcal{E},\Sigma(\mathcal{E}))$, shrinkage weights $\gamma$, and a bias vector $b\in\mathbb{R}^{p}$, let $\tilde{\omega}(\mathcal{E},\gamma)$ be as in (3). Define the two–sided, $(1-\eta)$ bias–aware lower and upper bounds for $\tau(\theta)$ by
$\ell_{b}(\mathcal{E},\Sigma,\gamma)$ $\equiv\hat{\tau}_{\gamma}\;-\;(\omega\circ(1-\gamma))^{\top}b\;-\;z_{1-\eta/2}\,\sqrt{\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)},$
$u_{b}(\mathcal{E},\Sigma,\gamma)$ $\equiv\hat{\tau}_{\gamma}\;-\;(\omega\circ(1-\gamma))^{\top}b\;+\;z_{1-\eta/2}\,\sqrt{\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)},$
@@
with $\alpha$ and $\beta$ as defined in Equation (4) and $\alpha^{\star},\beta^{\star}$ as defined in Equation (5).
@@
See Appendix C.1
Theorem 2 shows that the objective function is the same as the MSE-objective function up to a monotonic (square-root) transformation which does not affect the choice of the minimizer.
@@ 4.3 Vector valued estimands
Consider a vector-valued target estimand with entries $\tau^{\ell}(\theta)\equiv(\omega^{\ell})^{\top}\theta$ for $\ell=1,\dots L$, and $\omega^{\ell},\theta\in\mathbb{R}^{k}$.
For a subset $\mathcal{E}\subseteq\{1,\ldots,p\}$ and a positive-definite matrix $\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})\subset\mathbb{R}^{(p+|\mathcal{E}|)\times(p+|\mathcal{E}|)}$, define $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}^{p}$ an observational estimate and $\tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}\in\mathbb{R}^{|\mathcal{E}|}$ an experimental estimate each satisfying
for an unknown bias vector $b\in\mathbb{R}^{p}$. Moreover, define its joint variance as
$\mathbb{V}\!\begin{pmatrix}\tilde{\theta}^{\mathrm{obs}}-\theta\\[2.0pt] \tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}-\theta_{\mathcal{E}}\end{pmatrix}=\Sigma(\mathcal{E})$
assumed to be known, with uniformly bounded entries and strictly positive definite.
Given $(\mathcal{E},\Sigma)$, consider a class of linear plug-in estimators to allow the optimal shrinkage weights to depend on the specific entry of $\tau$.
@@
where $\gamma=(\gamma^{\ell}_{j})_{\ell\in[L],j\in\mathcal{E}}\in\mathbb{R}^{L\times|\mathcal{E}|}$ are shrinkage weights that can be an (implicit) function of $(\mathcal{E},\Sigma)$. Let $\hat{\theta}(\gamma)\equiv(\hat{\theta}^{1}(\gamma),\dots,\hat{\theta}^{L}(\gamma))\in\mathbb{R}^{p\times L}$ and $\hat{\tau}_{\gamma}\equiv(\hat{\tau}^{1}_{\gamma}(\hat{\theta}^{1}(\gamma)),\dots,\hat{\tau}^{L}_{\gamma}(\hat{\theta}^{L}(\gamma)))^{\top}\in\mathbb{R}^{L}$.
@@
Similar to the scalar case, we make the following assumption regarding the first-order estimation error.
@@
For any admissible $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$, assume for all $\ell=1,\dots,L$
@@
for a known weighting vector $\omega^{\ell}\in\mathbb{R}^{p}$, with $|\omega^{\ell}_{i}|>0$, for all $i,\ell$.
@@
As before, we introduce more compact notation before proceeding to the main result. For a given $(\mathcal{E},\gamma)$ define
@@
and let
$\tilde{\omega}^{\ell}(\mathcal{E},\gamma)\equiv\begin{pmatrix}\tilde{\omega}^{\ell}_{\text{obs}}(\mathcal{E},\gamma)\\[2.0pt] \tilde{\omega}^{\ell}_{\text{exp}}(\mathcal{E},\gamma)\end{pmatrix}\in\mathbb{R}^{\,p+|\mathcal{E}|}.$
Using this notation, we can write $\hat{\tau}^{\ell}_{\gamma}=\tilde{\omega}^{\ell}(\mathcal{E},\gamma)^{\top}\,\hat{\theta}$, where $\hat{\theta}=(\tilde{\theta}^{\text{obs}\top},\tilde{\theta}^{\text{exp}\top})^{\top}\in\mathbb{R}^{p+|\mathcal{E}|}$. Stacking across the $\ell$ index gives the estimator for the full vector $\tau$: $\hat{\tau}_{\gamma}=\tilde{\omega}(\mathcal{E},\gamma)\,\hat{\theta}$, where the $\ell^{th}$ row of $\tilde{\omega}(\mathcal{E},\gamma)\in\mathbb{R}^{L\times(p+|\mathcal{E}|)}$ is $\tilde{\omega}^{\ell}(\mathcal{E},\gamma)^{\top}$.
@@
where the dependence of $\bar{v}(\gamma)$ on $\mathcal{E}$ is omitted for conciseness.
The smallest worst-case variance and bias are defined as
@@ 5 Empirical applications
In this section we provide two empirical applications. In the first application we study the problem of choosing where to conduct an experiment using observational data from Banerjee et al. (2024). In the second application we illustrate the method for choosing treatment arms when estimating a structural model, calibrating our estimation to the PROGRESA experiment (Todd and Wolpin, 2006; Attanasio et al., 2012). We solve the optimization program by solving for each choice of experiments, both over $\gamma$ and the allocation of sample size via numerical optimizationand then enumerate the solutions across all possible choices.
5.1 Choosing experimental participants

### v2

@@ 5 Extensions
5.1 Local asymptotics in the presence of non-linearities
@@
This subsection extends our framework for (i) nonlinear moment functions $g(\theta)$ used in GMM in Section 4 and (ii) a smooth, potentially nonlinear estimand $\tau(\theta)$.
Linearization of the moments
Assume the sample moments obey root–$n$ scaling, i.e., $\mathbb{V}(\bar{g}_{\Sigma})=O(n^{-1})$. Let $b=b_{n}$ be a sequence where retain the same misspecification structure as in Section 4: the moment vector has mean $b=b_{n}$ with $(b_{n})_{j}=0$ for $j\in\mathcal{I}$ (experiment-based moments) and $(b_{n})_{\mathcal{I}^{c}}$ potentially nonzero. Our key assumption is a local condition on the moments bias of the form
$\|b_{n}\|^{2}\sqrt{n}\;\to\;0.$ (9)
This condition builds on the local asymptotic frameworks in Andrews et al. (2020) and Armstrong and Kolesár (2021). In their context, $b_{n}||^{2}=1/n$ matching the rate of convergence of the variance. In our case, $b_{n}$ can grow faster, slower or at the same rate of the standard error, encompassing $b_{n}||^{2}=1/n$ as a special case.
Under standard smoothness conditions, a GMM estimator with weighting matrix $W$ satisfies
$\sqrt{n}\big(\hat{\theta}(W,\Sigma)-\theta\big)=-(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W\,\sqrt{n}\big(\bar{g}_{\Sigma}-b_{n}\big)+o_{p}(1),$
so on the original (unscaled) scale the mean shift is $(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W\,b_{n}$ and the remainder is $o_{p}(n^{-1/2})$. The condition $\|b_{n}\|^{2}\sqrt{n}\to 0$ ensures that higher order terms are negligible under sufficient smoothness conditions, justifying the use of the linear expansion (5) for our design analysis. Because $b_{n}\to 0$, the Jacobian $\Lambda$ can be consistently estimated by evaluating the derivative of $g(\theta)$ at a preliminary observational estimator $\tilde{\theta}^{\mathrm{obs}}\rightarrow_{p}\theta$, assuming such estimates are available from an observational study (Andrews et al., 2020).
Linearization of the estimand.
@@
Let $\tau:\mathbb{R}^{p}\to\mathbb{R}$ be twice continuously differentiable near $\theta$, with gradient $\omega(\theta):=\partial\tau(\theta)/\partial\theta$ and bounded Hessian. A Taylor expansion around $\theta$ yields
$\tau(\hat{\theta})-\tau(\theta)=\omega(\theta)^{\top}(\hat{\theta}-\theta)\;+\;\tfrac{1}{2}\,(\hat{\theta}-\theta)^{\top}H_{\tau}(\bar{\theta})\,(\hat{\theta}-\theta),$
for some $\bar{\theta}$ between $\hat{\theta}$ and $\theta$, and $H_{\tau}$ is the Hessian. It follows that we can write
$\sqrt{n}\big(\tau(\hat{\theta})-\tau(\theta)\big)=\omega(\theta)^{\top}\,\Gamma_{\Lambda}(W)\,\sqrt{n}\big(\bar{g}_{\Sigma}-b_{n}\big)\;+\;o_{p}(1),$
because $\|\hat{\theta}-\theta\|=O_{p}(n^{-1/2}+\|b_{n}\|)$ and the local condition $\sqrt{n}\,\|b_{n}\|^{2}\to 0$ renders the quadratic remainder $o_{p}(1)$ after $\sqrt{n}$ scaling.
Consequently, the regret expressions derived for the linear case continue to apply up to $o_{p}(1)$, with $\omega$ interpreted as the gradient $\omega(\theta)$ (or any consistent plug-in $\omega(\tilde{\theta}^{\mathrm{obs}})$ from a preliminary estimator $\tilde{\theta}^{\mathrm{obs}}\to_{p}\theta$).
Remark 2 (Rescaling and change of variables linking Example 3).
An equivalent formulation uses $\sqrt{n}$–rescaled coordinates (as in Example 3). Define the rescaled moment vector $\tilde{g}:=\sqrt{n}\,\bar{g}$ and the rescaled estimation error $\tilde{\theta}\equiv\sqrt{n}\,(\hat{\theta}-\theta)$. Then
$\mathbb{V}(\tilde{g})=\Sigma=O(1),\qquad\mathbb{E}[\tilde{g}]=b_{n}\ \ \text{with}\ \ b_{n}\equiv\sqrt{n}\,b,$
so $b_{n}$ is the asymptotic bias and $\Sigma$ the asymptotic variance. The local misspecification condition in (9) is equivalent to $\frac{\|b_{n}\|^{2}}{\sqrt{n}}\;\to\;0,$ which coincides with Example 3. ∎
@@ 5.2 Confidence interval length as a regret criterion
In this subsection we consider an alternative loss based on the length of confidence intervals (CIs), focusing on Setting 1 for simplicity. For a candidate design $(\mathcal{E},\Sigma(\mathcal{E}))$, shrinkage weights $\gamma$, and a bias vector $b\in\mathbb{R}^{p}$, define the two–sided, $(1-\eta)$ lower and upper bounds for $\tau(\theta)$ by
$\ell_{b}(\mathcal{E},\Sigma,\gamma)$ $\equiv\hat{\tau}_{\gamma}\;-\;(\omega\circ(1-\gamma))^{\top}b\;-\;z_{1-\eta/2}\,\sqrt{\alpha(\mathcal{E},\Sigma,\gamma)},$
$u_{b}(\mathcal{E},\Sigma,\gamma)$ $\equiv\hat{\tau}_{\gamma}\;-\;(\omega\circ(1-\gamma))^{\top}b\;+\;z_{1-\eta/2}\,\sqrt{\alpha(\mathcal{E},\Sigma,\gamma)},$
@@
See Appendix B.1
Theorem 3 shows that the objective function is the same as the MSE-objective function up to a monotonic (square-root) transformation which does not affect the minimizer.
@@ 5.3 Vector valued estimands
Consider a vector-valued target estimand with entries $\tau^{\ell}(\theta)\equiv(\omega^{\ell})^{\top}\theta$ for $\ell=1,\dots q$, and $\omega^{\ell},\theta\in\mathbb{R}^{k}$.
Given $(\mathcal{E},\Sigma)$, and observational and experimental estimates $\tilde{\theta}^{\mathrm{obs}},\tilde{\theta}^{\mathrm{exp}}$, consider a class of linear plug-in estimators to allow the optimal shrinkage weights to depend on the specific entry of $\tau$.
@@
where $\gamma=(\gamma^{\ell}_{j})_{\ell\in\{1,\cdots,q\},j\in\mathcal{E}}\in\mathbb{R}^{q\times|\mathcal{E}|}$ are shrinkage weights that can be an (implicit) function of $(\mathcal{E},\Sigma)$. Let $\hat{\theta}(\gamma)\equiv(\hat{\theta}^{1}(\gamma),\dots,\hat{\theta}^{q}(\gamma))\in\mathbb{R}^{p\times q}$ and $\hat{\tau}_{\gamma}\equiv(\hat{\tau}^{1}_{\gamma}(\hat{\theta}^{1}(\gamma)),\dots,\hat{\tau}^{q}_{\gamma}(\hat{\theta}^{q}(\gamma)))^{\top}\in\mathbb{R}^{q}$.
@@
For any admissible $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$, assume for all $\ell=1,\dots,q$
@@
for a known weighting vector $\omega^{\ell}\in\mathbb{R}^{p}$, with $|\omega^{\ell}_{i}|\in(0,\infty)$, for all $i,\ell$.
@@
We introduce more compact notation before proceeding to the main result. For a given $(\mathcal{E},\gamma)$ define
@@
and let $\tilde{\omega}^{\ell}(\mathcal{E},\gamma)\equiv\begin{pmatrix}\tilde{\omega}^{\ell}_{\text{obs}}(\mathcal{E},\gamma)\\[2.0pt] \tilde{\omega}^{\ell}_{\text{exp}}(\mathcal{E},\gamma)\end{pmatrix}\in\mathbb{R}^{\,p+|\mathcal{E}|}.$ We can write $\hat{\tau}^{\ell}_{\gamma}=\tilde{\omega}^{\ell}(\mathcal{E},\gamma)^{\top}\,\hat{\theta}$, where $\hat{\theta}=(\tilde{\theta}^{\text{obs}\top},\tilde{\theta}^{\text{exp}\top})^{\top}\in\mathbb{R}^{p+|\mathcal{E}|}$. Stacking across the $\ell$ index gives the estimator for the full vector $\tau$: $\hat{\tau}_{\gamma}=\tilde{\omega}(\mathcal{E},\gamma)\,\hat{\theta}$, where the $\ell^{th}$ row of $\tilde{\omega}(\mathcal{E},\gamma)\in\mathbb{R}^{q\times(p+|\mathcal{E}|)}$ is $\tilde{\omega}^{\ell}(\mathcal{E},\gamma)^{\top}$.
@@
where the dependence of $\bar{v}(\gamma)$ on $\mathcal{E}$ is omitted for conciseness and $\circ$ denotes the component-wise multiplication. The smallest worst-case variance and bias are defined as
@@ 6 Empirical applications
In this section we provide two empirical applications, one for measuring GE effect and the second one for externally valid site selection.
6.1 An implementation guide for measuring GE effects
Combining experimental or pre-program observational data with structural models is increasingly used for program evaluation (Todd and Wolpin, 2006; Attanasio et al., 2012; Meghir et al., 2022; Allcott et al., 2025). [^8] Across all AEA articles in 2015-2025, about $10\%$ of the experimental papers combine experimental results with structural models (i.e., one-third of the papers that use observational in combination with experimental estimates). Because large-scale experimentation are often infeasible due to cost constraints (Muralidharan and Niehaus, 2017) these models are meant to extrapolate effects from small-scale experiments to learn general-equilibrium (GE) effects. This extrapolation typically relies on other experiments or observational data.
We present a (step-by-step) illustration below when these models are calibrated using previous studies conducted in a different country.
Question
Consider a researcher evaluating a cash transfers for sending children to school in rural Kenya where such a program is not in place. Due to budget constraints, the researcher can only randomize at a small scale (partial equilibrium), but ultimately wishes to predict GE effects. Researchers have access to preliminary estimates from the Mexican PROGRESA experiment (Todd and Wolpin, 2006), which we may be concerned, lack external validity.
6.1.1 Step 1: model description
The first step for a researcher is to describe the modeling choices. We present a stylized model of school choice from Bonhomme and Weidner (2022) and Todd and Wolpin (2006).
Individual choices
Let $S\in\{0,1\}$ denote school attendance, $C$ consumption, $Y$ (pre-transfer) household income, $W$ the child’s potential wage, and $t$ the stipend when enrolled. Abstracting from covariates (we will introduce covariates in the estimation), utility is
with budget C=Y+W(1S)+tSC=Y+W(1-S)+tS. The parametrization $(\xi_{3}-\xi_{1}-\xi_{2})$ is without loss and simplifies expressions below. Enrollment satisfies $S=1\big\{U(Y+t,1,t,\varepsilon)>U(Y+W,0,0,0)\big\}.$ Letting $Z(Y,W,t)\equiv\xi_{1}W-\xi_{2}Y-\xi_{3}t-\xi_{4},$ we have $P(S=1\mid Y,W,t)=\Phi\!\big(-Z(Y,W,t)\big)$, with $\Phi(\cdot)$ denoting the standard Gaussian CDF.
General equilibrium effects
We are interested in the effect of a small stipend $t$ to all eligible (poor) households in rural Kenya. GE feedback is allowed through income and wages: for functions $y(t),w(t)$ and mean-zero idiosyncratic income and wage shocks $\varepsilon_{Yi},\varepsilon_{Wi}$, let
which decomposes into the direct stipend effect $\phi_{0}\xi_{3}$ and the indirect GE effects via income and wages, $\phi_{0}\xi_{2}y_{0}$ and $\phi_{0}\xi_{1}w_{0}$. Under simple market clearing for labor supply and demand, we can show [^9] This follows from the following: set $L_{s}(W,t)=L_{d}(W)$, the labor supply equals the labor demand, and differentiate at $t=0$: $\frac{\partial L_{s}}{\partial W}w_{0}+\frac{\partial L_{s}}{\partial t}=\frac{\partial L_{d}}{\partial W}w_{0}$. Suppose we have a constant number of hours worked for working child, denoted as $H$, and any child not at school is working. From the probit index, $\partial_{t}\mathbb{E}[S\,|\,\cdot]=\phi(-Z)\,\xi_{3}$ and $\partial_{W}\mathbb{E}[S\,|\,\cdot]=-\phi(-Z)\,\xi_{1}$ at $t=0$. Hence $\partial L_{s}/\partial t=-\phi_{0}\xi_{3}H$ and $\partial L_{s}/\partial W=\phi_{0}\xi_{1}H$. Dividing the equation by $H$ we obtain the desired result.
$w_{0}=\;\frac{\phi_{0}\,\xi_{3}}{\phi_{0}\,\xi_{1}-d},$ (15)
where $d$ denotes the slope of the demand curve divided by total hours of work.
6.1.2 Step 2: parametrization and transparency on biases
Given the model, the following step for researchers is to define the parameters of interest. Here, we define the parameters of interest as
with corresponding estimand
$\tau(\theta)\;=\;\underbrace{\theta_{2}}_{\text{Direct effect}}\;+\;\underbrace{y_{0}}_{\text{Income multiplier}}\,\underbrace{\theta_{1}}_{\text{Income effect}}\;+\;\underbrace{w_{0}(\theta)\,\theta_{3}}_{\text{Wage effect}},\qquad w_{0}(\theta)\;=\;\frac{\theta_{2}}{-\theta_{3}-d},$ (17)
where the explicit form of $w_{0}(\theta)$ follows from (15).
We treat $d$ as known for simplicity, using meta-analyses on demand elasiticities, and $y_{0}=1.5$ because computed experimentally in Kenya by Egger et al. (2022). [^10] Egger et al. (2022) report a total income/consumption multiplier $M_{\text{tot}}=\partial_{t}\mathbb{E}\big[Y_{\text{pre}}(t)+t\big]|_{t=0}=2.5$; therefore the target derivative of pre-transfer income is $y_{0}=\partial_{t}\mathbb{E}\big[Y_{\text{pre}}(t)\big]|_{t=0}=M_{\text{tot}}-1$. Specifically, $d=0.5\frac{1-S_{0}}{W_{0}}$, where $S_{0}$ and $W_{0}$ are baseline probability to go to school (assuming all children not in school go to work) and baseline wages calibrated to the pre-experimental data from PROGRESA and $0.5$ obtained from meta-studies in Espey and Thilmany (2000). [^11] One could also calibrate $S_{0},W_{0}$ to pre-experimental data in Kenya directly when available.
In this example, the primary concern is bias in the observational estimator $\tilde{\theta}^{\mathrm{obs}}$ constructed as described below, abstracting from potential biases in $y_{0},d$ which, for simplicity, we assume to be second-order. When these biases are considered first-order, researchers should let $y_{0},d$ be additional parameters in $\theta$. This difference illustrates a feature of our framework: the chosen parameterization ensures transparency about which biases’ concerns drive the design; these considerations can be collected in a pre-analysis plan.
6.1.3 Step 3: observational study estimates
Using experimental data from Mexico’s PROGRESA program, we estimate $(\xi_{1},\xi_{2},\xi_{3},\xi_{4})$ via probit with standard controls (age, distance to school, eligibility, year, highest grade attained). Estimation is conducted separately by gender, and we will be focusing here on the effects on female students.
We pool treated and control observations to estimate the schooling effect $\xi_{4}$, controlling for the individual-level subsidy. In turn, we obtain observational estimates $\big(\tilde{\theta}_{1}^{\mathrm{obs}},\tilde{\theta}_{2}^{\mathrm{obs}},\tilde{\theta}_{3}^{\mathrm{obs}}\big)$ that map to the model parametrization. Standard errors are constructed using the Delta method with clustering at the village level. Following Section 5.1, we then construct $\omega=\frac{\partial\tau}{\partial\theta}\Big|_{\theta=\tilde{\theta}^{\mathrm{obs}}}$. Point estimates, estimated sensitivity $\omega$, and standard errors are reported below.
Parameter $\tilde{\theta}^{\mathrm{obs}}$ $\omega$ $\sigma$
[…]
We observe heterogeneity in both the sensitivity weights $\omega$ the precision of the observational estimates.
6.1.4 Step 4: describing the candidate experiments
The following step is to list the feasible experiments, subject to budget constraints.
In this example, we consider a researcher who can afford only small, partial-equilibrium experiments in Kenya, and examine three possible experiments for a given sample size $n$:
    with precision $v_{1}^{2}/n$. The researcher optimizes over $\gamma_{1}$, while $\theta_{2}$ and $\theta_{3}$ are calibrated to the PROGRESA study.
    with precision $v_{2}^{2}/n$. The researcher optimizes over $\gamma_{2}$, while $\theta_{1}$ and $\theta_{3}$ are calibrated to the PROGRESA study.
For simplicity, we calibrate $v_{1}^{2}$ as $\sigma_{1}^{2}\times n^{\mathrm{obs}}$, where $\sigma_{1}^{2}$ is the variance of $\tilde{\theta}_{1}^{\mathrm{obs}}$ and $n^{\mathrm{obs}}$ is the observational-sample size ($n^{\mathrm{obs}}=1089$ for the female sample using data from eligible students in Todd and Wolpin (2006)). We calibrate $v_{2}^{2}$ analogously as $\sigma_{2}^{2}\times n^{\mathrm{obs}}$. In practice, researchers may use pilot studies to calibrate the experimental variances when available.
Experiment type Identified parameter Choice Variables Calibrated standard error Description
6.1.5 Step 5: experimental design
Figure 4: Regret–optimal sample allocation across treatment arms when both the unconditional cash transfer (UCT; income shock) and the conditional cash transfer (CCT; stipend) arms are feasible. The optimal solution is interior: the vast majority of participants (over 90%) are assigned to CCT, with a small but nonzero fraction assigned to UCT to hedge misspecification. This pattern reflects the higher payoff to learning about CCT in this application relative to UCT.
The final step is to design the experiment and report it in a pre-analysis plan.
Choice of the experiment when only one treatment arm is available
Figure 5 (right panel) displays the selected arm as a function of $n$ (together with the corresponding $\gamma_{j}^{\star}$). We encode “not chosen” as $\gamma_{j}^{\star}=0$. For small samples ($n\lesssim 500$), the optimal choice is UCT: the experimental CCT estimator would be too noisy, so the resulting variance under CCT dominates, while the bias advantage of CCT over UCT is limited because $\omega_{1}$ and $\omega_{2}$ are of similar magnitude. Once $n$ crosses $n\approx 500$, CCT becomes dominant: by combining observational and experimental estimates for CCT, we can achieve comparable variance, and the worst-case bias is lower than under UCT.
Choice of the estimator
In addition to sample allocation, our method optimizes the shrinkage weights of the parameters of interest. Figure 5 (left panel) shows that when both arms can be run, we place essentially full weight on the experimental estimator for UCT ($\gamma^{\star}_{\text{UCT}}\approx 1$ across $n$), while for CCT we gradually shift toward the experimental estimator as precision improves: $\gamma^{\star}_{\text{CCT}}$ rises from about $0.5$ at small $n$ to approximately $0.9$ by $n=1000$. Intuitively, higher $n$ lowers experimental variance, making it optimal to rely more heavily on experimental evidence for the CCT parameter.
Sample size allocation when both arms are feasible
Figure 4 reports the sample allocation implied by the regret-optimal design when both treatment arms are feasible. The optimal allocation is interior: most participants (over roughly $90\%$) are assigned to the CCT arm (education), with a small but nonzero fraction assigned to the UCT arm (income shock). Two forces rationalize this pattern. First, the CCT parameter is more misspecification-sensitive ($\omega_{2}\approx 1.98$ vs. $\omega_{1}\approx 1.5$), so generating experimental evidence on CCT has a larger payoff in bias reduction. Second, the income effect $\theta_{1}$’s variance is about an order of magnitude smaller than the stipend effect $\theta_{2}$’s variance in our PROGRESA calibration, lowering the marginal returns to learn about UCT relative to CCT. Hence, while the designer keeps some allocation on UCT, most of the sample is placed on CCT; our framework makes this bias–variance trade-off explicit.
Regret comparisons
Figure 6 reports worst-case regret for each design choice. The left panel shows that the regret of our method (green line, running both arms with an interior allocation) declines toward one as $n$ increases, reflecting near-oracle performance under our normalization. In contrast, single-arm designs exhibit nonvanishing regret because bias remains first order even as variance shrinks. The right panel focuses on the single-arm problem: both the oracle and the researcher can only choose a single arm. In this case, UCT is preferable only for small $n$; beyond the same threshold ($\approx 500$), CCT yields uniformly lower regret as variance becomes second order relative to bias. The regret vanishes relative to the oracle that can only choose a single arm.
Implications
When only one arm is feasible, CCT is preferable only once the total sample size is sufficiently large; otherwise UCT is preferred. This pattern reflects the interaction between bias and variance: for small samples, CCT’s bias advantage is not large enough to compensate for its higher variance relative to UCT, whereas at larger samples the bias reduction becomes dominant. Existing designs based on variance considerations would fail to provide solutions that balance bias and variance comparisons.
When feasible, it is instead optimal to run two arms with highly unbalanced sample, reflecting the relative contribution of the variance for each treatment arm.
Following these steps, researchers can similarly report their design in a pre-analysis plan.
6.2 Measuring the performance of our method for site selection